# Neural Population Geometry and Optimal Coding of Tasks with Shared Latent Structure

### 26 Feb 2024 v1

@@ 3.3 Geometry of Multi-task Learning in MLPs
Figure 5: Theory predicts generalization error of the Hebb rule in trained and random MLPs. (a) Schematic of the experiment. Latent variables $z$ are randomly shattered to generate task labels. These latents are passed through a random MLP (light blue) and are then used as inputs to train a 3 hidden layer MLP (dark blue) on the multi-task binary classification problem using stochastic gradient descent. (b-c) After training, we sample a new set of latents and teacher vectors and calculate the generalization error of the Hebb rule on each layer of the (b) random and (c) trained network. Theoretical predictions closely track empirical errors, and the trained network achieves a lower error in intermediate layers. (d-f) Geometric terms across layers for the random (light blue line) and trained (dark blue line) networks. Linear layers are marked by circles and relu layers by squares. Interestingly, the error only slightly changes across linear and relu layers of the same model stage, in spite of sharp changes in the geometry. (g-i) Average change in the geometry between relu and linear layers. Here we show the mean and standard deviation of the difference between each geometric term before and after the relu nonlinearity is applied. The application of relu consistently causes increases in the dimension and alignment, as well as decreases in the correlation.
To test our theory on explicitly non-Gaussian data, we now apply our theory to both random and trained non-linear multi-layer perceptrons (MLPs). As shown in Fig. 5a, we first sample a set of zero-mean Gaussian latent variables $z$. We then feed these latent variables through a random MLP with increasing hidden layer sizes. After generating this set of stimuli, we sample $500$ random $T$ vectors, and use these to generate a set of task labels (Methods). These labels and data points from the random MLP are then used to train a downstream 4-hidden-layer MLP, which predicts the label for each task using a linear readout of its penultimate layer. This is reminiscent of the hidden manifold modeling framework studied in the machine learning community [30]. Finally, we sample a new set of latent variables along with a new set of tasks and calculate the generalization error of the Hebb rule when applied to the representations at different layers of the trained and untrained MLPs (Methods). This setup allows us to validate our theory on non-linear transformations of Gaussian latent variables.
@@
Turning to our geometric decomposition of the error, we find several interesting trends across layers and non-linearities (Fig. 5d-f). In the random MLP, we see that the correlation decreases throughout the network, while the alignment term exhibits small fluctuations while gradually decreasing overall. Since we choose the layer sizes to be monotonically increasing through the network, we also find that the dimension of the responses stays more or less constant through the network.
@@
In the trained network, we find that linear and relu layers orchestrate a trade off between dimensionality, alignment, and correlation that together only slightly perturb the generalization error. As shown in Fig 5g-i, the neural dimensionality and alignment spike each time the nonlinearity is applied, at the cost of the total correlation. Conversely, at each linear layer, the correlation sharply rises while the dimensionality and alignment fall. Importantly, this pattern is not present in the random network. Thus, the trained MLP learns to use the nonlinearity to increase the dimension of the representations while simultaneously squashing particularly harmful noisy directions of variability in the input and increasing how well factorized the code is. Intuitively, the linear layers learn to orient signal-unrelated features into parts of the state space that are zeroed out by the non-linearity [37]. While these are sharp changes in the geometry, the overall generalization error across layers exhibits only minor fluctuations between most linear and relu layers, highlighting the limitations of a generalization error-based approach to analyzing network activity.
@@
We now study the dynamics of these geometric terms through training. To do this, we track the layer-wise generalization error and geometry through a single epoch of training (Methods). As shown in Fig. 6, we find that the generalization error in this context decays through training across layers. In the linear layers, we find a small increase in the total correlation and dimension as well as a steady rise in the alignment. On the other hand, in relu layers, improved generalization is driven by an increase in the dimension and alignment across both early and late layers. In these layers, the total correlation decreases throughout training. We find similar trends for MLPs with a $\tanh$ non-linearity (SM). These geometric trends mirror the effects of adding more samples on the optimal representation: As more samples become available, the alignment and dimension grow while the correlation shrinks. In other words, the dynamics of MLP representations in non-linear layers qualitatively mimic the behavior of the optimal neural code described by our theory.
@@
Figure 6: Evolution of generalization error and representational geometry through training. (a) Generalization error of the Hebb rule applied to four different layers on a previously unseen set of tasks. We examine layers from the early stages of the MLP (black) as well as late stages (green) for both linear (solid line) and relu (dashed line) layers. We can see that the error decreases across all layers with the number of training samples, $p$. (b) Dimension term. While the dimension of the neural activations increases for all layers through training, the differences are most pronounced in the relu layers. (c) Correlation term. This term decreases through training for relu layers and slightly increases for linear layers. (d) Alignment term. This uniformly increases through training across the network. These findings mirror the effect of increasing the number of training samples on the optimal neural code.
@@ 3.4 Predicting Readout Performance of Macaque Visual Representations
We find that our formula accurately predicts the generalization error of the monkey neural responses across both recording sites. Furthermore, as shown in Fig. 7d, both V4 and IT achieve better generalization error than the raw pixels, and IT achieves a better generalization error than V4. [38]. Applying our geometric decomposition of the generalization error to these data, we find that the dimensionality is lower in V4 than in either IT or the pixel data [22] (Fig. 7e). Furthermore, we find that the alignment is far lower in V4 than in IT, suggesting that factorization of these latent variables increases along the ventral stream, in line with previous results [39]. However, we find that the alignment is higher in the pixels than in either V4 or IT. This finding suggests that the ventral stream representations considered here do not factorize information regarding object position well, as a poor alignment indicates that neuronal noise and the coding directions of distinct latents are not orthogonal to one another (Fig. 7e-g). Nevertheless, we do find that the total correlation is much higher in V4 and IT than in the pixel space. This suggests that while the representations considered here are not well factorized, they do contain correlations with the latent variables that can be leveraged to achieve a generalization error that is better than classification on the pixels. Taken together, these results provide strong support for our theory and demonstrate the applicability of our metrics as a tool for tying the geometry of neuronal population responses to the computational objective of multi-task learning.
Figure 7: Theory predicts multi-task error in macaque V4 and IT data. (a-b) Example stimuli and tasks. Visual stimuli included 64 objects grouped into 8 categories and were generated by modifying $d=6$ continuous latent variables that included object size, position, and angle. We form binary classification tasks on subsets of the data coming from the same category–e.g. all images from the ”Tables” category. (c) Probe placement. Figure adapted with permission from [33]. (d) Generalization error across pixels and brain regions calculated empirically (markers) and using our formula for the generalization error (solid line). (e-g) Geometric terms across pixels and neural responses. We can see that V4 representations are lower dimensional than either the pixels or IT, and that both IT and V4 suffer from a lower alignment than the pixels. However, both regions present with a much higher total correlation with the latent variables, enabling improved generalization performance.
@@ 4 Discussion
In this work, we analyzed a model of learning multiple tasks that share a common latent structure, connecting neural population geometry to the multi-task learning problem. In this model, binary classification tasks are generated by shattering a latent space uniformly. By varying the spectrum of the latent covariance, we were able to model situations in which certain latent variables are, on average, more useful for downstream tasks than others. We then calculated the generalization error of a linear readout applied to a set of neural activations responding to stimuli that were generated from these latent variables. In this way, we evaluated how well a neural representation is able to extract information about latent variables from the stimuli and use this for arbitrary downstream tasks.
@@
To connect the mesoscopic statistics of neural activation patterns to generalization performance, we decomposed our formula for the generalization error into three geometric terms. These terms summarize the relationship between population activity and the latent variables and completely determine generalization performance in our task. First is the dimensionality of the entire population activity, as measured by the participation ratio. Second, we introduced the total correlation between neuronal activity and the latent variables, representing the overall sensitivity of neuronal responses to changes in the latent space. Finally, we defined the alignment, which summarizes the degree to which distinct latent variables and latent-unrelated variability are coded along orthogonal directions in the neural state space. As such, this term measures the degree to which a representation is factorized and manages to keep noise variability out of the coding directions. Together, these three terms completely characterize the generalization performance of the linear readout applied to the neural population responses.
@@
To investigate the role that latent structure and data scarcity play in shaping population codes, we next analytically calculated which geometries minimize the generalization error, given a fixed latent structure and number of available training samples. This falls in line with studies on efficient coding in which predictions of sensory physiology are derived from statistics of the natural environment [40, 41, 42]. In this way, we found that disentangled (or factorized) representations naturally emerge as an optimal solution to the multi-task learning problem. Specifically, we found that the eigenvectors of the neuronal and latent covariances match the singular vectors of the cross-covariance (Methods). This is reminiscent of work done on the successor representation and grid cells [43, 44, 45, 46, 47], and it would be interesting to see if there are additional constraints under which grid cells or other coding properties emerge as an optimal code in our model [48].
@@
We then described how the optimal neuronal code shifts from being low to high dimensional as the number of available samples increases. This finding reflects a strategy of compressing less useful information when data is scarce and expanding it when data is abundant. It is possible that a similar relation holds for low vs. high noise learning scenarios [49, 50, 51, 52]. It is interesting to compare these findings to previous work [22] reporting that higher dimensional representations of individual object classes are preferred for invariant object recognition in the few shot learning setting. Although the contribution to the error of the dimensionality scales as 1/ p in our formula just as in this work, the additional contribution to the error of the alignment term, leads to optimal representations with lower dimensionality in the few shot regime. These results highlight the fact that different geometric terms may compete with each other in ways that cannot be directly read off from generalization error equations when there are additional constraints imposed on the system. Finally, we note that theories of least squares estimation suggest that similar phenomena occur for more complex readouts [53, 54, 55, 56]. We suspect that the same spectral expansion trend we describe holds for such readouts, though perhaps the optimal code can remain smooth in such cases [27]. It would be interesting to investigate these questions either analytically or numerically in future work.
@@
To test our theory, we next considered a setting in which a set of Gaussian latents are fed through a random network, and task labels are generated as above in the latent space. A four hidden layer MLP was then trained on the outputs of the random network on these tasks [30]. We then evaluated each layer of the trained network on previously unseen tasks using the scheme described above. Here, we found that intermediate layers of the trained network successfully learned to represent the latent variables in a way that could be used for downstream tasks, in line with previous work [7]. Furthermore, we found a strong correspondence between our theoretical predictions and the empirical generalization errors. Applying our geometric decomposition of the error, we found that linear and relu layers obtained similar generalization errors but with sharply different underlying geometries. Specifically, relu layers yielded a large boost to dimensionality and alignment at the cost of the total correlation. This suggests a strategy in which linear layers attempt to place neuronal noise in regions of the state space that are then squashed by the non-linearity [37]. Tracking the evolution of these terms through training, we found that the correlation in relu layers decreases through training, while the alignment and dimension increase. This mirrors the behavior of an optimal code as more samples are presented as well as the behavior of deep linear networks trained on certain regression tasks [57, 58]. It would be interesting to apply our geometric decomposition to analyze disentanglement in more complex artificial networks.
@@
Finally, we applied our theory to electrophysiological recordings from macaque V4 and IT [33]. Specifically, we used our theory to predict the generalization error of a linear readout of neural responses trained on a series of binary classification tasks. We formed task labels using latent variables describing object position, size, and orientation. As before, our theory showed excellent agreement with empirical generalization errors of linear readouts applied to the neural responses. We found that both regions showed better readout performance than the pixels themselves [38]. While we found a relatively low alignment in these neural responses, we did find that alignment improved from V4 to IT. These findings suggest that, while these latent variables are not overall well factorized in the ventral stream, the degree of factorization improves from V4 to IT, partially in line with previous results [39]. It would be very interesting to repeat this experiment with visual regions specialized for localization. Overall, these results demonstrate the applicability of our theory as a data-analytic tool.
While we primarily focused on neural representations, it would be very interesting to derive behavioral predictions from this theory. Specifically, one could decompose the generalization error formula into error rates for easy vs. hard examples. In this way, we could leverage our theory to predict learning curve dynamics across stimuli of varying levels of difficulty. These predictions could then be tested in perceptual discrimination tasks involving continuous modulation of an unobserved latent space. We hope to pursue this line of research in subsequent work.

### 11 Apr 2024 v2

@@ 3.3 Geometry of multi-task learning in MLPs
Figure 5: Theory predicts generalization error of the Hebb rule in trained and random MLPs. (a) Schematic of the experiment. Latent variables $z$ are randomly shattered to generate task labels. These latents are passed through a random MLP (light blue) and are then used as inputs to train a 3 hidden layer MLP (dark blue) on the multi-task binary classification problem using stochastic gradient descent. (b-c) After training, we sample a new set of latents and teacher vectors and calculate the generalization error of the Hebb rule on each layer of the (b) random and (c) trained network. Theoretical predictions closely track empirical errors, and the trained network achieves a lower error in intermediate layers. (d-g) Geometric terms across layers for the random (light blue line) and trained (dark blue line) networks. Linear layers are marked by circles and relu layers by squares. Interestingly, the error only slightly changes across linear and relu layers of the same model stage, in spite of sharp changes in the geometry. (h-k) Average change in the geometry before and after the application of relu. Here we show the mean and standard deviation of the difference between each geometric term before and after the relu nonlinearity is applied. In the trained network, the application of relu consistently causes increases in the dimension and signal-signal factorization, as well as decreases in the correlation.
In practice, neural data is complex and may be non-Gaussian. To test our theory in this regime, we now apply our theory to both random and trained non-linear multi-layer perceptrons (MLPs). As shown in Fig. 5a, we first sample a set of zero-mean Gaussian latent variables $z$. We then feed these latent variables through a random MLP with increasing hidden layer sizes. After generating this set of stimuli, we sample $500$ random $T$ vectors, and use these to generate a set of task labels (Methods). These labels and data points from the random MLP are then used to train a downstream 4-hidden-layer MLP, which predicts the label for each task using a linear readout of its penultimate layer. This is reminiscent of the hidden manifold modeling framework [36]. Finally, we sample a new set of latent variables along with a new set of tasks and calculate the generalization error of the Hebb rule when applied to the representations at different layers of the trained and untrained MLPs (Methods). This setup allows us to validate our theory on non-linear transformations of Gaussian latent variables.
@@
Turning to our geometric decomposition of the error, we find several interesting trends across layers and non-linearities (Fig. 5d-g). In the random MLP, we see that the correlation decreases each time the non-linearity is applied, while the signal-noise and signal-signal factorization terms gradually decrease through the network. We also find that the dimension of the responses stays more or less constant through the network.
@@
In the trained network, we find that linear and relu layers orchestrate a trade off between the geometric terms that together lead to an overall decrease in multi-task generalization error. As shown in Fig 5h-k, the neural dimension spikes each time the nonlinearity is applied, at the cost of the total correlation. Conversely, at each linear layer, the correlation sharply rises while the dimension falls. On the other hand, we find that both the signal-signal and signal-noise factorization terms monotonically increase through the penultimate layer of the network. Importantly, this pattern is not present in the random network. Thus, the trained MLP learns to use the nonlinearity to increase the dimension of the representation as well as the factorization of the latent variables, while simultaneously squashing particularly harmful noisy directions of variability in the input. Intuitively, the linear layers learn to orient signal-unrelated features into parts of the state space that are zeroed out by the non-linearity [43]. While these are sharp changes in the geometry, the overall generalization error across layers exhibits only minor fluctuations between most linear and relu layers, highlighting the limitations of a generalization error-based approach to analyzing network activity.
@@
We now study the dynamics of these geometric terms through training. To do this, we track the layer-wise generalization error and geometry through a single epoch of training (Methods). As shown in Fig. 6, we find that the generalization error in this context decays through training across layers. In the linear layers, we find that improved generalization is driven by a steady rise in the signal-signal and signal-noise factorization as well as small increases in the total correlation and dimension. On the other hand, in relu layers, the correlation decreases through training, with improved generalization being driven by increases in the dimension as well as the signal-signal and signal-noise factorizations across both early and late layers. We find similar trends for MLPs with a $\tanh$ non-linearity (SM Figs. 2&3). These geometric trends mirror the effects of adding more samples on the optimal representation: As more samples become available, the factorization terms and dimension grow while the correlation shrinks. In other words, the dynamics of MLP representations in non-linear layers qualitatively mimic the behavior of the optimal neural code described by our theory.
@@
Figure 6: Evolution of generalization error and representational geometry through training. (a) Generalization error of the Hebb rule applied to four different layers on a previously unseen set of tasks. We examine layers from the early stages of the MLP (black) as well as late stages (green) for both linear (solid line) and relu (dashed line) layers. We can see that the error decreases across all layers with the number of training samples, $p$. (b) Dimension term. While the dimension of the neural activations increases for all layers through training, the differences are most pronounced in the relu layers. (c) Correlation term. This term decreases through training for relu layers and slightly increases for linear layers. (d) Signal-signal factorization term. (e) Signal-noise factorization term. Both of these terms uniformly increase through training across the network, though the former begins to plateau at the end of training. These findings mirror the effect of increasing the number of training samples on the optimal neural code.
@@ 3.4 Predicting readout performance of macaque visual representations
We find that our formula accurately predicts the generalization error of the monkey neural responses across both recording sites. Furthermore, as shown in Fig. 7d, both V4 and IT achieve better generalization error than the raw pixels, and IT achieves a better generalization error than V4. [44]. Applying our geometric decomposition of the generalization error to these data, we find that the dimension is lower in V4 than in either IT or the pixel data [26] (Fig. 7e). Furthermore, we find that the correlation steadily increases from the raw pixels to V4 and IT, indicating an increased overall signal. We also find that the level of signal-signal factorization improves from V4 to IT, in line with previous results [45]. Interestingly, we find that the signal-signal factorization is higher in the pixel space than in either brain region. Finally, we find that the signal-noise factorization is far lower in the pixels and V4 than in IT. This suggests that in IT, latent-unrelated variability overlaps less with the coding directions than in V4 or the pixels (Fig. 7e-h). Taken together, these results provide strong support for our theory and demonstrate the applicability of our metrics as a tool for tying the geometry of neuronal population responses to the computational objective of multi-task learning.
Figure 7: Theory predicts multi-task error in macaque V4 and IT data. (a-b) Example stimuli and tasks. Visual stimuli included 64 objects grouped into 8 categories and were generated by modifying $d=6$ continuous latent variables that included object size, position, and angle. We form binary classification tasks on subsets of the data coming from the same category–e.g. all images from the ”Tables” category. (c) Probe placement. Figure adapted with permission from [39]. (d) Generalization error across pixels and brain regions calculated empirically (markers) and using our formula for the generalization error (solid line). (e-h) Geometric terms across pixels and neural responses. We can see that V4 representations are lower dimensional than either the pixels or IT, that the correlation improves throughout, and that the signal-signal factorization is highest in the pixel space, though it improves from V4 to IT. Note too that the signal-noise factorization sharply rises from V4 to IT, suggesting that signal-unrelated variability overlaps less with the coding directions in the latter region.
@@ 4 Discussion
In this work, we analyzed a model of learning multiple tasks that share a common latent structure, connecting neural population geometry to the multi-task learning problem. In this model, binary classification tasks are generated by shattering a latent space uniformly. By varying the spectrum of the latent covariance matrix, we were able to model situations in which certain latent variables are, on average, more useful for downstream tasks than others. We then calculated the generalization error of a linear readout applied to a set of neural activations responding to stimuli that were generated from these latent variables. In this way, we evaluated how well a neural representation is able to extract information about latent variables from the stimuli and use this for arbitrary downstream tasks.
@@
To connect the mesoscopic statistics of neural activation patterns to generalization performance, we decomposed our formula for the generalization error into four geometric terms. These terms summarize the relationship between population activity and the latent variables and completely determine generalization performance in our task. First is the dimension of the entire population activity, as measured by the participation ratio [46]. Second, we introduced the total correlation between neuronal activity and the latent variables, representing the overall sensitivity of neuronal responses to changes in the latent space. Third, we introduced the signal-signal factorization, measuring the degree to which each latent variable is coded in an orthogonal direction in the neural state space, with adequate variance devoted to each coding direction. Finally, we defined the signal-noise factorization, which summarizes the amount of signal-unrelated variability that lies along the directions in state space used to represent the latent variables. Hence, this term summarizes the degree to which noise correlations affect the readout considered here [47, 48, 49, 50]. Together, these four terms completely characterize the generalization performance of the linear readout applied to the neural population responses.
@@
We next analytically calculated which geometries minimize the generalization error, given a fixed latent structure and number of available training samples. This falls in line with studies on efficient coding in which predictions of sensory physiology are derived from statistics of the natural environment [51, 52, 53]. In this way, we found that disentangled representations naturally emerge as an optimal solution to the multi-task learning problem. Specifically, we found that the eigenvectors of the neuronal and latent covariances match the singular vectors of the cross-covariance (Methods, SM). This is reminiscent of work done on the successor representation and grid cells [54, 55, 56, 57, 58], and it would be interesting to see if there are additional constraints under which grid cells or other coding properties emerge as an optimal code in our model [59]. Future work can also extend our theory to consider sets of tasks that compete with one another as in previous studies of multi-task learning [60, 61]. Moreover, given recent findings regarding multi-task learning and dynamical motif reuse in recurrent networks [62, 63], it would be interesting to extend our theory to a dynamical setting in which predictions are formed using an appropriate readout of time-varying firing rates.
We then described how the optimal neuronal code shifts from being low to high dimensional as the number of available samples increases. This finding reflects a strategy of compressing less useful information when data is scarce and expanding it when data is abundant. It would be very interesting to test this normative prediction in behaving animals. This could be done, for example, using multi-unit recordings taken while an animal learns to perform decision making tasks in which multiple latent variables (e.g. distinct context cues) must be used to make a decision. Our theory predicts that task-related neural variability in the early stages of training should primarily track highly informative latent variables. As training progresses, we predict that task-relevant neural variability will increase in dimension as it begins to track the less informative latent variables in the task.
@@
It is interesting to compare these findings to previous work reporting that higher dimensional representations of individual object classes are preferred for invariant object recognition in the few shot learning setting [26]. Although the contribution to the error of the neural dimension scales as 1/ p in our formula just as in this work, the additional contribution to the error of the signal-signal and signal-noise factorization terms together with the constraint of of the covariance matrix remaining positive semi-definite, leads to optimal representations with lower dimension in the few shot regime. These results highlight the fact that different geometric terms may compete with each other in ways that cannot be directly read off from generalization error equations when there are additional constraints imposed on the system. Another related line of work has considered the role of neural dimension in random pattern separation [64, 65, 66]. In these settings, the optimal dimension can decrease when neural noise increases, and future work can examine whether a similar relation holds in the multi-task learning setting considered here. Finally, we note that theories of least squares estimation suggest that similar phenomena occur for regression tasks with more complex readouts [67, 68, 69, 70]. We suspect that a similar spectral expansion trend of optimal coding schemes holds for other common readout mechanisms in both classification and regression tasks, and it would be interesting to investigate these questions in future work.
@@
To test our theory, we next considered a setting in which a set of Gaussian latents are fed through a random network, and task labels are generated as above in the latent space. A four hidden layer MLP was then trained on the outputs of the random network on these tasks [36]. We then evaluated each layer of the trained network on previously unseen tasks using the scheme described above. Here, we found that intermediate layers of the trained network successfully learned to represent the latent variables in a way that could be used for downstream tasks, in line with previous work [8]. Furthermore, we found a strong correspondence between our theoretical predictions and the empirical generalization errors. Applying our geometric decomposition of the error, we found that linear and relu layers obtained similar generalization errors but with sharply different underlying geometries. Specifically, relu layers yielded a large boost to the dimension at the cost of the total correlation, while the signal-signal and signal-noise factorizations steadily improved through the network. This suggests a strategy in which linear layers attempt to place neuronal noise in regions of the state space that are then squashed by the non-linearity [43]. Tracking the evolution of these terms through training, we found that the correlation in relu layers decreases through training, while the dimension, signal-signal factorization, and signal-noise factorization increase. This mirrors the behavior of an optimal code as more samples are presented as well as the behavior of deep linear networks trained on certain regression tasks [71, 72]. It would be interesting to apply our geometric decomposition to analyze disentanglement in more complex artificial networks.
@@
Finally, we applied our theory to electrophysiological recordings from macaque V4 and IT [39]. Specifically, we used our theory to predict the generalization error of a linear readout of neural responses trained on a series of binary classification tasks. We formed task labels using latent variables describing object position, size, and orientation. As before, our theory showed excellent agreement with empirical generalization errors of linear readouts applied to the neural responses. We found that both regions showed better readout performance than the pixels themselves [44]. While we found a low signal-signal factorization in these neural responses relative to the pixels, we did find that signal-signal factorization improved from V4 to IT [45]. Moreover, we found a sharp increase in the signal-noise factorization from the pixels and V4 to IT, suggesting that latent-unrelated variability may become increasingly orthogonal to the coding directions through the ventral stream. It would be very interesting to repeat such an analysis using data from visual regions specialized for localization. Overall, these results demonstrate the applicability of our theory as a data-analytic tool.
While we focused on tasks that came from shattering a continuous latent space, future work could extend this theory to consider distributions of latents and tasks that more closely mirror common experimental settings. Specifically, one could repeat our general calculation, while restricting the distribution of task vectors to a handful of relevant directions or restricting the distribution of latent variables to be fixed to discrete values. This would be particularly interesting in settings where only particular groups of tasks or latent variable values are relevant for an animal to consider. Such a calculation would allow one to tie the population geometry to the linear readout performance on families of tasks and latent distributions which are tailored to specific experimental settings. We hope to pursue this line of research in subsequent work.